OpenStudy (anonymous):

–12 ÷ 3 • (–8 + (–4)2 – 6) +2

OpenStudy (anonymous):

it's not 4)2 it's 4 to the 2nd power

hartnn (hartnn):

so whats \((-4)^2\) = ?

OpenStudy (anonymous):

16?

hartnn (hartnn):

yup,sice u solve parenthesis first whats (-8+16-6) = ?

hartnn (hartnn):

yup,sice u solve parenthesis first whats (-8+16-6) = ?

OpenStudy (anonymous):

–12 ÷ 3 • (–8 + (–4)2 – 6) +2 -12 divided by 3 * (-8 + 16 - 6) + 2 -12 divided by 3 * (+2) + 2 -12 divided by 3 * 2 + 2 -4 * 2 + 2 -2 + 2 0

OpenStudy (anonymous):

It's 2 right barton?

hartnn (hartnn):

yup :) which gives u -12/3*2+2 use PEDMAS, and whats -12/3 ?

OpenStudy (anonymous):

-6?

OpenStudy (anonymous):

-4 :)

hartnn (hartnn):

yup :) so now u have -4*2+2 again PEDMAS requires u to do -4*2 first so whats -4*2 ??

OpenStudy (anonymous):

-8

hartnn (hartnn):

yup,last step -8+2 ?

OpenStudy (anonymous):

-6

hartnn (hartnn):

thats your final answer. :)

OpenStudy (anonymous):

You. Are. Awesome!

hartnn (hartnn):

thank you. u are awesome too :)